OpenStudy (anonymous):

solve the system of equations using any method? 3x+2y=11 x-2=-4y

OpenStudy (anonymous):

matrices bro

OpenStudy (waleed_imtiaz):

3x+2y=11.... multiply it with 2 to equal the co-efficients of y..... NOw subtract the two equations..... U will get 6x+4y=22 -(x+4y=2) =5x+0y=20 ...... U will get x=4... put x in any of the equations. U will get y=-1/2

OpenStudy (anonymous):

can you show the work @waleed_imtiaz

OpenStudy (waleed_imtiaz):

Yes...... Offcourse..... I said Multiply the first equation with 2. OKay... U will get 6x+4y=22..... now subtract this multiplied equation with the second equation x+4y=2... NOw U can subtract it urself... After subtraction. the y component will be vanished..... and u will get 5x=20.... divide 5 on both sides..... x=4... U got it........? and let me put x=4 in second equation..... x+4y=2.. 4y=2-4=-2... y=-2/4=-1/2...... any confusion ?